OpenStudy (anonymous):

Show the expanded form and simplification of x^4/x^2 Explain in your own words how you can simplify x^4/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (campbell_st):

use the index law for division \[\frac{x^a}{x^b} = x^{a - b}\] you have \[\frac{x^4}{x^2} = \] same base subtract the powers..